An object is launched from ground level directly upward at a rate of 48 meters per second. The equation for the object’s height is y = -16x2 + 48x. What values of x is the object at or above a height of 32 meters?
Answer by ikleyn(52800)

.

To answer this question, solve equation

    -16x2 + 48x = 32.


Cancel the common factor 16 to simplify

    -x^2 + 3x = 2

    x^2 - 3x + 2 = 0

    (x-1)*(x-2) = 0.


The roots are  x= 1  and  x= 2.


So, an object will be at or above the height 0f 32 meters  from x= 1 second to x= 2 seconds.    ANSWER

Solved.

------------------

In this site,  there is a bunch of lessons on a projectile thrown/shot/launched vertically up
    - Introductory lesson on a projectile thrown-shot-launched vertically up
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.
Read them attentively and learn how to solve this type of problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.